# UFO: A Unified Approach to Fine-grained Visual Perception via Open-ended Language Interface

### v2

@@ 4 Training
Table 2: Results on GiT [61]’s multi-task benchmark. “single-task” refers to models trained on each task separately, while “multi-task” indicates models trained jointly across all selected tasks. $\star$” denotes the model is capable of the task, though no number is reported. “-” means incapability in that specific task. We highlight the top-1 entries of one-stage multi-task generalist models and joint training improvements with bold font. We follow [61] to list specific modules, which are designed for specific functions in architecture.
@@
Methods Examples Num #Params AP AP50 AP75 AP AP50 AP75 mIoU(SS) BLEU-4 CIDEr Acc@0.5
@@
Improvement${}_{~{}\text{(single}\rightarrow\text{multi)}}$ +0.5 +0.9 +0.6 +0.9 +0.4 +0.9 +0.7 +1.1 +3.1 +2.2
@@
Task Example Sources Count Size
Visual Question Answering LLaVA-v1.5- mix665k [39] 1 0.7m
Refer Expression Comprehension  RefCOCO [75], RefCLEF [28] 4 0.8m
Refer Expression Segmentation  RefCOCO [75], RefCLEF [28] 4 0.8m
Object Detection Objects365 [56], COCO [37] 5 3.8m
Instance Segmentation OpenImages [32], LVIS [23] 6 1.4m
Semantic Segmentation COCOStuff [3], ADE20K [80] 4 0.3m
Table 3: Datasets used in instruction tuning. The columns, from left to right, represent task name, dataset examples, dataset count, and data amount. More details are provided in the appendix.
@@ 4.1 Multi-Task Training
Architecture. To ensure fair comparison and validate the effectiveness of our approach across various architectures, we conduct multi-task training using two variants: UFO-ViT and UFO-InternVL2-8B. UFO-ViT strictly follows GiT [61], employing a SAM [30]-pretrained ViT [21] and a text tokenizer from BERT [19]. It is available in three sizes: ViT-B, ViT-L, and ViT-H. UFO-InternVL2-8B utilizes the pretraining weight of InternVL2-8B [12], with detailed model specifications provided in Table 1.
Datasets. We use the same multi-task dataset as GiT, which includes five standard visual perception tasks: COCO 2017 [37] for object detection and instance segmentation, COCO Caption [10] for image captioning, the RefCOCO series [43, 75] for referring expression comprehension (REC), and ADE20K [80] for semantic segmentation.
@@ 4.2 Fine-grained Instruction Tuning
Architecture. To demonstrate that our method is applicable to various MLLMs, we use not only InternVL2-8B [12] but also the LLaVA-1.5-7B [40] for pretraining, specifically UFO-LLaVA-1.5-7B. Architecture details are in Table 1.
Datasets. To enhance the model’s versatility, we enrich our training datasets with more datasets and tasks, extensively covering major visual perception tasks Our training data includes 24 datasets across 6 tasks, as detailed in Table 3. More details of data composition are in the appendix.
@@
LISA [33] - - - - - - - - - 74.1 76.5 71.1 62.4 67.4 56.5 66.4 68.5 67.9
PixelLM [54] - - - - - - - - - 73.0 76.5 68.2 66.3 71.7 58.3 69.3 70.5 69.2
SAM4MLLM-7B [11] - - - - - - - - - 77.1 80.9 72.5 71.5 76.8 64.7 74.5 75.2 74.2
NExT-Chat [76] 85.5 90.0 77.9 77.2 84.5 68.0 80.1 79.8 80.4 74.7 78.9 69.5 65.1 71.9 56.7 67.0 67.0 68.9
PerceptionGPT [47] 88.6 92.5 84.6 82.1 88.6 74.2 84.1 85.2 85.0 75.1 78.6 71.7 68.5 73.9 61.3 70.3 71.7 71.4
@@
Ferret-7B [72] 87.5 91.4 82.5 80.8 87.4 73.1 83.9 84.8 83.9 - - - - - - - - -
VistaLLM [48] 88.1 91.5 83.0 82.9 89.8 74.8 83.6 84.4 84.8 74.5 76.0 72.7 69.1 73.7 64.0 69.0 70.9 71.2
UFO-LLaVA-1.5-7B 89.9 93.0 86.0 84.3 90.3 77.5 86.0 86.8 86.7 76.2 78.5 73.0 69.8 75.1 63.9 71.1 71.9 72.4
UFO-LLaVA-1.5-7B* 90.8 93.0 87.3 85.5 90.5 78.6 86.9 87.2 87.5 77.2 79.4 73.8 70.8 75.5 64.1 72.1 73.2 73.3
UFO-InternVL2-8B 90.7 94.0 87.4 85.4 90.1 79.2 86.7 87.0 87.6 77.3 79.2 74.7 71.3 75.7 66.5 72.8 73.2 73.8
UFO-InternVL2-8B* 91.4 93.8 88.2 85.7 90.7 79.7 86.8 87.4 88.0 78.0 79.7 75.6 72.3 76.8 66.6 73.7 74.3 74.6
5 Experiments
5.1 Experimental Settings
Multi-Task Training Details. To facilitate comparison and verify the effectiveness of our method, we also conduct single-task training independently on five selected tasks. For both single-task and multi-task training, we use a batch size of 24 and employ the AdamW [29] optimizer with a cosine annealing schedule, setting the initial learning rate to 0.0002. More details are in the appendix.
Fine-grained Instruction Tuning Details. In training, we use a batch size of 32 with gradient accumulation set to 16, running on 8 NVIDIA A100 GPUs for 120K iterations. The AdamW [29] optimizer and a cosine annealing schedule are employed, with a learning rate of 0.0002 and weight decay of 0.01. For efficient training, we employ LoRA [25] with a rank of 8, freezing the image tokenizer while keeping only the LLM trainable. More details are in the appendix.
Training Objectives. All tasks utilize a CrossEntropy Loss as they are unified under the open-ended language interface. For segmentation tasks, we additionally apply focal loss [55] and dice loss to supervise the binary mask output. The final loss for segmentation tasks is expressed as:
We empirically find that setting all weights to 1 offers better overall performance. See appendix for more details.
@@ 5.1 Experimental Settings
LISA-LLaVA-7B [33] 36.8 37.6 36.6
LISA-LLaVA-7B [33]* 47.3 40.6 49.4
LISA-LLaVA-1.5-7B [33] 48.7 47.1 49.2
LISA-LLaVA-1.5-7B [33]* 55.6 48.3 57.9
UFO-LLaVA-1.5-7B 53.8 40.1 58.2
UFO-LLaVA-1.5-7B* 58.0 46.3 61.7
UFO-InternVL2-8B 55.4 41.9 59.8
UFO-InternVL2-8B* 61.2 49.6 64.9
Table 5: Results (gIoU) on ReasonSeg test set. * denotes using 239 reasoning segmentation data samples to fine-tune the model.
5.2 Multi-Task Evaluation
We evaluate our model’s performance in both single-task and multi-task settings across five standard vision-centric tasks, benchmarking it against specialized and generalist models. Without extra task decoders, our model adapts to various tasks through the proposed open-ended language interface and achieves outstanding performance.
Comparison with Specialist Models. As shown in Table 2, our single-task model effectively bridges the performance gap with specialized models, achieving either superior or on-par performance. For example, we achieve 47.8 mAP in detection compared to 45.4 mAP with Deformable-DETR [83] and 49.5 mIoU in semantic segmentation against 47.2 mIoU with Mask2Former [13]. In instance segmentation, we also outperform specialized methods like Mask R-CNN [24] and Polar Mask [70] while matching the performance of Mask2Former [13].
Comparison with Generalist Models. To facilitate comparison with GiT [61], we adopt its one-stage training method without task-specific finetuning. This involves jointly training on a mixed dataset of the five tasks and directly testing on their respective validation or test sets. Table 2 shows that our model outperforms the previous leading generalist model, GiT, across all tasks, with the same pretraining and data. Notably, in the largest model size, we outperform GiT by 12.3 mAP on COCO instance segmentation and 3.3 mIoU on ADE20K semantic segmentation, demonstrating the superiority of our segmentation modeling. The improvement in captioning is mainly because we use a standard vocabulary for all tasks, while GiT uses task-specific vocabularies, hindering the learning of general text representation.
After scaling to MLLMs, we observe improved performance on captioning and REC, while performance on other tasks remains comparable to the UFO-ViT-L. We speculate that this performance difference primarily arises from different pretraining. For UFO-ViT, we use SAM [30] pretraining, making it more aligned with detection and segmentation. In contrast, InternVL2-8B is mainly pre-trained on image-level vision-language tasks, which better suit captioning and REC.
5.3 Fine-grained Instruction Tuning Results
After training MLLMs with data shown in Table 3, we first evaluate their object and pixel-level perception performance on the visual grounding tasks. Then we evaluate on a more difficult task, reasoning segmentation, examining whether our method effectively combines fine-grained perception with the MLLMs’ language reasoning abilities.
Visual Grounding maps textual descriptions to objects in an image, which can be further categorized into referring expression comprehension (REC) and segmentation (RES). We comprehensively list the results for the two tasks in Table 4. We report results in two settings: direct evaluation after joint training and specifically finetuning. Without using box decoders, our best model surpasses the previous state-of-the-art VisionLLM v2 [68] by an average of 2.0%. After specific finetuning, our model achieves stronger performance with 88.0% on average. While all previous approaches rely on mask decoders or polygon approximations for segmentation, our method delivers superior or comparable performance without them. For instance, our InternVL2 variant outperforms the SAM [30]-based LISA [33] by an average of 5.9 cIoU and matches with VisionLLM v2 [68]. These outcomes validate the effectiveness of our method, demonstrating that with proper task modeling, MLLMs can handle fine-grained perception tasks without task decoders.
Reasoning Segmentation (ReasonSeg) is a challenging benchmark introduced by LISA [33], which presents more sophisticated and nuanced instructions, requiring models to leverage world knowledge and engage in deeper logical reasoning. We adhere to the evaluation protocol of LISA, reporting both zero-shot and finetuned results. As shown in Table 5, with the same pretraining, our LLaVA-1.5 variant outperforms LISA by 5.1 gIoU in zero-shot settings. With advanced MLLM InternVL2, our performance further improves, exceeding LISA by 6.7 gIoU. Finetuning also significantly improves performance, with our InternVL2 variant improving by 5.8 gIoU compared to zero-shot.
Since reasoning segmentation requires both language reasoning and precise segmentation, we attribute our improvement to better task integration through unified modeling. In LISA, the MLLM handles only language reasoning and generates coarse segmentation prompts, relying on an additional SAM for finer segmentation. This leads to information loss and insufficient synergy. In our unified modeling, the MLLM manages both language reasoning and precise segmentation, allowing different task capabilities to fully integrate within a shared parameter space, thereby enhancing synergy.

### v3

@@ 4 Training
Table 2: Results on GiT [69]’s multi-task benchmark. $\star$” denotes the model is capable of the task but no number is reported. “-” means incapability. We highlight joint training improvements with bold font and follow [69] to list modules for specific functions.
@@
Methods Examples Num #Params AP AP50 AP75 AP AP50 AP75 mIoU($S$) BLEU-4 CIDEr Acc@0.5
@@
[…]
@@ 4.1 Multi-Task Training
Architecture. To ensure fair comparison and validate our effectiveness across various architectures, we conduct multi-task training using two variants: UFO-ViT and UFO-InternVL2.5-8B. UFO-ViT strictly follows GiT [69], employing a SAM [34]-pretrained ViT [21] and a text tokenizer from BERT [19]. It is available in three sizes: ViT-B, ViT-L, and ViT-H. UFO-InternVL2.5-8B utilizes the pretraining weight of InternVL2.5-8B [12], with detailed model specifications provided in Table 1.
Datasets. We use the same multi-task dataset as GiT: COCO 2017 [42] for object detection and instance segmentation, COCO Caption [10] for image captioning, the RefCOCO series [48, 83] for referring expression comprehension (REC), and ADE20K [90] for semantic segmentation.
@@ 4.2 Fine-grained Instruction Tuning
Architecture. To demonstrate that our method is applicable to various MLLMs, we use not only InternVL2.5-8B [12] but also the LLaVA-1.5-7B [44] for pretraining, specifically UFO-LLaVA-1.5-7B. Architecture details are in Table 1.
Datasets. To enhance the model’s versatility, we enrich the training data to 2.5M across 6 tasks, including VQA data from [39], COCO-Stuff [4], LVIS [25], etc. We additionally add RES task on the basis of five tasks in multi-task training. More details of data composition are in Table 9.
5 Experiments
5.1 Experimental Settings
Multi-Task Training Details. To facilitate comparison with specialist models, we also conduct single-task training independently on five selected tasks. For both single-task and multi-task training, we use a batch size of 24 and employ the AdamW [33] optimizer with a cosine annealing schedule, setting the initial learning rate to 0.0002. More details are in the appendix.
Fine-grained Instruction Tuning Details. In training, we use a batch size of 32 with gradient accumulation set to 16, running on 8 NVIDIA A100 GPUs for 120K iterations. The AdamW [33] optimizer and a cosine annealing schedule are employed, with a learning rate of 0.0002 and weight decay of 0.01. For efficient training, we employ LoRA [27] with a rank of 8, freezing the image tokenizer while keeping only the LLM trainable. More details are in Appendix Table 12.
Training Objectives. All tasks utilize a CrossEntropy Loss as they are unified under the open-ended language interface. For segmentation tasks, we additionally apply focal loss [61] and dice loss to supervise the mask output. The final loss for segmentation tasks is expressed as:
We find that setting all weights to 1 offers better overall performance. See appendix for more details.
@@ 5.1 Experimental Settings
GLaMM-7B [57] - - - - - - - - - 79.5 83.2 76.9 72.6 78.7 64.6 74.2 74.9 75.6
SAM4MLLM-8B [11] - - - - - - - - - 79.8 82.7 74.7 74.6 80.0 67.2 75.5 76.4 76.4
HiMTok-8B [73] - - - - - - - - - 81.1 81.2 79.2 77.1 78.8 71.5 75.8 76.7 77.7
PerceptionGPT-7B [53] 88.6 92.5 84.6 82.1 88.6 74.2 84.1 85.2 85.0 75.1 78.6 71.7 68.5 73.9 61.3 70.3 71.7 71.4
@@
Ferret-v2-7B [85] 92.8 94.7 88.7 87.4 92.8 79.3 89.4 89.3 89.3 - - - - - - - - -
VistaLLM-7B [54] 88.1 91.5 83.0 82.9 89.8 74.8 83.6 84.4 84.8 74.5 76.0 72.7 69.1 73.7 64.0 69.0 70.9 71.2
UFO-LLaVA-1.5-7B 90.2 93.5 87.3 84.4 90.3 78.7 86.4 86.8 87.2 77.2 80.1 76.4 71.8 77.9 70.2 74.1 73.5 75.2
UFO-LLaVA-1.5-7B* 91.1 93.7 88.6 85.5 90.5 79.9 87.3 87.2 88.0 77.9 81.1 77.0 72.5 78.5 71.4 75.6 74.1 76.0
UFO-InternVL2.5-8B 91.8 94.3 87.5 86.9 91.3 80.6 87.9 88.6 88.6 80.0 81.6 78.1 76.7 79.9 72.3 75.5 76.3 77.6
UFO-InternVL2.5-8B* 93.1 94.8 89.2 87.7 92.1 82.3 88.2 89.2 89.6 81.0 82.6 78.6 77.1 80.4 72.6 76.7 77.3 78.3
Table 4: Results (gIoU) on ReasonSeg test set. * using reasoning segmentation data in training.
@@
LISA-7B [37] 36.8 37.6 36.6
LISA-7B [37]* 47.3 40.6 49.4
Cores-7B [3] 48.7 41.0 50.9
Cores-7B [3]* 52.4 44.2 55.0
HiMTok-8B [73]* 60.8 - -
UFO-LLaVA-1.5-7B 54.4 41.2 58.5
UFO-LLaVA-1.5-7B* 58.8 46.5 62.7
UFO-InternVL2.5-8B 60.0 48.7 63.6
UFO-InternVL2.5-8B* 67.0 56.2 70.4
Table 5: Results on vision-language benchmarks.
Models GQA MMBench MMVP HallBench
InternVL2.5-8B 60.6 84.6 76.3 50.1
UFO-InternVL2.5-8B 60.8 84.2 76.3 50.7
[…]
$N^{2}$ 1 4 16 25
mAP 38.9 41.3 42.6 42.9
FPS 7.0 5.9 3.6 2.8
[…]